# Interpreting the Second-Order Effects of Neurons in CLIP

### v2

@@ 4 Sparse decomposition of neurons
We aim to interpret each neuron by associating its second-order effect with text. We build on the previous observation that each second-order effect of a neuron $\phi_{n}^{l}$ is associated with a vector direction $r_{n}^{l}$. Since $r_{n}^{l}$ lies in a shared image-text space, we can decompose it to a sparse set of text directions. We use a sparse coding method [30] to mine for a small set of texts for each neuron, out of a large pool of descriptions. We evaluate the found texts across different initial pools with different set sizes.
@@
Table 1: Examples of sparse decompositions. For each neuron, we present the top-4 texts corresponding to the sparse decomposition with $m=128$ and the signs of the coefficients in the decomposition for two initial pools.
@@
Figure 6: Images with largest second-order effect norm per neuron. We present the top images from 10% of ImageNet test set for the neurons in Tab. 1. Note that neurons are polysemantic - they have large second-order effects on images that show multiple concepts (e.g. cars and boats).
Experimental settings. We verify that the reconstructed $\hat{r}^{l}_{n}$ from the text representations captures the variation in the image representation, as measured by zero-shot accuracy on IN. We simultaneously replace the neurons’ second-order contributions in a single layer with the approximation $x^{l}_{n}(I)\hat{r}^{l}_{n}+b^{l}_{n}$.
To obtain sparse decomposition for each neuron, we use scikit-learn’s implementation of orthogonal matching pursuit [30]. We consider two strategies for constructing the pool of text descriptions $T$. The first type is single words - the 10k and 30k most common words in English. The second type is image descriptions - we prompt ChatGPT-3.5 to produce descriptions of images that include an object of a specific class. Repeating this process for all the IN classes results in $\sim$28k unique image descriptions. We then evaluate the reconstruction of $r^{l}_{n}$ for different $m$’s and pools.
Effect of sparse set size $m$ and different pools. We experiment with $m\in\{4,8,16,32,64,128\}$ and the three text pools, and present the accuracy on 10% of IN test set in Fig. 5. We approach the original classification accuracy with 128 text descriptions per neuron reconstruction $\hat{r}^{l}_{n}$. Using full descriptions outperforms using single words for the text pool, but the gap vanishes for larger $m$.
Qualitative results. We present the images with the largest second-order norms in Fig. 6, and the corresponding top-4 text descriptions in Tab. 1. As shown, the found descriptions match the objects in the top 10 images. Moreover, some individual neurons correspond to multiple concepts (e.g. writing both toward “yacht” and a type of a car - “cabriolet”). This corroborates with previous literature on neurons’ polysemantic behavior [11] - single neurons behave as a superposition of multiple interpretable features. This property will allow us to generate adversarial images in Sec. 5.1.
@@ 5.1 Automatic generation of adversarial examples
The sparse decomposition of $r_{n}^{l}$’s allows us to find overlapping concepts that neurons are writing to. We use these spurious cues to generate semantic adversarial images. Our pipeline, shown in Fig. 1, mines for spurious words that correlate with the incorrect class in CLIP (e.g. “elephant”, that correlates with “dog”), combines them into image descriptions that include the correct class name (“cat”), and generates adversarial images by providing these descriptions to a text-to-image model. We explain the steps in the pipeline and provide quantitative and qualitative results.
Figure 7: Adversarial images generated by our method. For each binary classification task, we present the generated images, the input text to the text-to-image model (words from $W^{v}$ are bold), and an attribution map [13] for the classification (areas that contribute to the incorrect class score are red).
@@
Experimental settings. We generate adversarial images for classifying between pairs of classes from CIFAR-10 [21]. We use the 30k most common words as our pool $T$. We choose the top $100$ neurons from layers 8-10 for $\mathcal{N}$, and the top 25 words according to their contribution scores for prompting the LLM. We prompt LLaMA3 [43] to generate 50 descriptions for each classification task (see prompt in App. A.6). We then filter out descriptions that include the class name and choose 10 random descriptions. We generate 10 images for each description with DeepFloyd IF text-to-image model [42]. This results in 100 images per experiment. We repeat the experiment 3 times and manually remove images that include $c_{2}$ objects or do not include $c_{1}$ objects.
We report three additional baselines. First, we repeat the same process with $100$ random neurons instead of the set $\mathcal{N}$. Second, we repeat the same generation process with sparse text decompositions computed from the first principle components of the indirect effects instead of the second-order effect. Third, we do not rely on the neuron decompositions, and instead prompt the language model with the words from $M$ for which their text representations are the most similar to $v$. Both for our pipeline and the baselines, we automatically filter out synonyms of $c_{2}$ from the phrases provided to the language model according to their sentence similarity to $c_{2}$ [35].
Quantitative results. The classification accuracy results for the adversarial images are presented in Tab. 3. The success rate of our adversarial images is significantly higher than the indirect effect baseline, the similar words baseline, and the random baseline, which succeeds only accidentally.
Qualitative results. Fig. 7 includes generated adversarial examples and the descriptions that were used in their generation. The presented attribution heatmaps [13] show that the found spurious objects from $W_{v}$ contribute the most to the misclassification, while the object from the correct class (e.g. a horse in the left-most image) contributes the least. We provide more results for additional classification tasks (e.g. “stop-sign v.s. yield”) in Fig. 14.
@@
[…]
Figure 9: Qualitative results on IN-Segmentation (ViT-B-32). Our heatmaps capture more object parts than the first-order token decomposition of [13].
5.2 Concept discovery in images
We aim to discover concepts in image $I$, by aggregating phrases that correspond to the neurons that are activated on $I$. Here, we start from the set of activated neurons $\mathcal{N}$ (for which $||\phi_{n}^{l}(I)||_{2}$ is above the 98th percentile of norms computed across IN images). Similarly to the contribution score described earlier, we compute an image-contribution score $w_{j}^{I}$ for each phrase $j$ according to its combined weight in the decompositions of neurons in $\mathcal{N}$. Formally, $w_{j}^{I}$ is the overall sum of weights that each neuron in $\mathcal{N}$ assigns to $j$ in its decomposition, weighted by the neuron second-order norms: $w_{j}^{I}=\sum_{n\in{\mathcal{N}}}{\gamma_{j}^{l,n}||\phi_{n}^{l}(I)||_{2}}$. The phrases with the highest image-contribution score are picked to describe the image concepts.
Qualitative results. We present qualitative results for neurons and the top-10 discovered concepts from layer 9 of ViT-B-32 in Fig. 9, when using the most common words as the pool. The number of neurons activated on these images, $|\mathcal{N}|$, is between 29 and 59, less than 2% of the neurons in the layer. Nevertheless, the top words extracted from these neurons relate semantically to the objects in the image and their locations. Surprisingly, the top word for each of the images appears only in one or two of the neuron sparse decompositions and is not spread across many activated neurons.
[…]
@@ 5.2 Concept discovery in images
horse $\rightarrow$ 1.0 ($\pm$1.4) 2.8 ($\pm$3.7) 1.0 ($\pm$1.4) 5.3 ($\pm$1.9)
automobile
@@
ship $\rightarrow$ 1.3 ($\pm$1.9) 0.0 ($\pm$0.0) 1.3 ($\pm$0.9) 7.0 ($\pm$4.5)
automobile
@@
Partial-LRP [44] 55.0 35.5 66.9
Rollout [1] 61.8 42.6 74.0
LRP [6] 62.9 35.8 68.5
GradCAM [39] 67.3 39.3 61.9
Chefer et al. [8] 68.9 49.1 79.7
@@
Table 3: Segmentation performance on IN-segmentation. Our zero-shot segmentation is more accurate than previous methods across all metrics.
@@ 5.3 Zero-shot segmentation
Specifically, to generate a segmentation map for an image $I$, we find a set of neurons with the largest absolute value of the dot product with the encoded class name we aim to segment: $|\langle r_{n}^{l},M_{\text{text}}(c_{i})\rangle|$. We then average their spatial activation maps $p_{i}^{l,n}(I)$, standardize the average activations into $[0,1]$, and binarize the values into foreground/background segments by applying a threshold of 0.5.
Segmentation results. We provide results on IN-Segmentation [16], which includes foreground/background segmentation maps of IN objects. We use activation maps from the top 200 neurons of layers 8-10. Tab. 3 presents a quantitative comparison to previous explainability methods. Our method outperforms other zero-shot segmentation methods across all standard evaluation metrics. We provide qualitative results before thresholding in Fig. 9. While the first-order effects highlight individual discriminative object parts, our heatmaps capture more parts of the full object.
@@ 6 Limitations and discussion
We analyzed the second-order effects of neurons on the CLIP representation and used our understanding to perform zero-shot segmentation, discover image concepts, and generate adversarial images. We present mechanisms that we did not analyze in our investigation and conclude with a discussion of broader impact and future directions.
@@
Future work and broader impact. The mass production of adversarial images can be harmful to systems that rely on neural networks. On the other hand, automatic extraction of such cases allows the defender to be prepared for them and, possibly, fine-tune the model on the generated images to avoid such attacks. We plan to investigate this approach to improve CLIP’s robustness in the future.
@@ A.1 Second order ablations for ViT-L
We repeat the same experiments from Sec. 3.3 for ViT-L-14, trained on LAION dataset [38]. For this model, we only use 10% of ImageNet test set. Here, the maximal drop in performance when ablating the second order is relatively smaller and is spread across more layers. Nevertheless, the same properties presented and discussed in Section 3.3 hold for this model.
@@ A.2 First order ablations
For the two models discussed above, ViT-B-32 and ViT-L-14, we provide the mean-ablation results for the first-order effects of MSA layers, as computed in [13]. For each model, we present the performance before and after accumulative mean-ablation of all the first-order effects of MSA layers. As shown in Figures 12 and 12, the neurons with the significant second-order effects appear right before the layers with the significant first-order effects.
@@ A.4 Additional sparse decomposition results
We provide additional examples of sparse decompositions of neurons in Tab. 4 and the images with the top norms for the second-order effects of the same neurons in Fig. 13. As shown, the found descriptions match the objects in the top 10 images.
A.5 Derivations with Layer Normalization
@@
where $x\in\mathbb{R}^{d^{\prime}}$ is the input token, $\mu_{l},\sigma_{l}\in\mathbb{R}$ are the mean and standard deviation, and $\gamma,\beta\in\mathbb{R}^{d^{\prime}}$ are learned vectors. To include $A$ and $B$ in the second-order effect of a neuron flow, we replace the input-independent component in Eq. 5, $PW_{VO}^{l^{\prime},h}w^{l^{\prime},n}$, with:
@@
where $LN^{l}$ is the normalization layer at layer $l$, that can be parameterized similarly to Eq. 8 by $A^{l},B^{l}\in\mathbb{R}^{d^{\prime}}$. We modify the definition of the second-order effect accordingly:
@@
Figure 13: Images with largest second-order effect norm per neuron. We present the top images from 10% of ImageNet test set, for the neurons in Tab. 4.
@@
A.6 Prompts
We provide the prompt that was used for generating sentences given the set of words $W_{v}$, as presented in Section 5.1, in Tab. 5. This prompt is given to LLAMA3 model [43].
Additionally, we provide the prompt that was used for generating the pool of ImageNet class descriptions, presented in Section 4. We prompt ChatGPT (GPT 3.5) with the prompt template provided in Tab. 6.
A.7 Compute

### v3

@@ 4 Sparse decomposition of neurons
We aim to interpret each neuron by associating its second-order effect with text. We build on the previous observation that each second-order effect of a neuron $\phi_{n}^{l}$ is associated with a vector direction $r_{n}^{l}$. Since $r_{n}^{l}$ lies in a shared image-text space, we can decompose it to a sparse set of text directions. We use a sparse coding method (Pati et al., 1993) to mine for a small set of texts for each neuron, out of a large pool of descriptions. We evaluate the found texts across different initial pools with different set sizes.
@@
Table 1: Examples of sparse decompositions (ViT-B-32, layer 9). We present the top-4 texts corresponding to the sparse decomposition of each neuron and the signs of the decomposition coefficients, for two initial pools ($m=128$). See Table 4 for more neurons.
@@
Figure 6: Images with largest second-order effect norm per neuron. We present the top images from 10% of ImageNet validation set for the neurons in Table 1. Note that neurons are polysemantic - they have large second-order effects on images that show multiple concepts (e.g. cars and boats). See top-50 images in Figure 14.
Experimental settings. We verify that the reconstructed $\hat{r}^{l}_{n}$ from the text representations captures the variation in the image representation, as measured by zero-shot accuracy on ImageNet. We simultaneously replace the neurons’ second-order contributions in a single layer with the approximation $x^{l}_{n}(I)\hat{r}^{l}_{n}+b^{l}_{n}$.
To obtain sparse decomposition for each neuron, we use scikit-learn’s implementation of orthogonal matching pursuit (Pati et al., 1993). We consider two strategies for constructing the pool of text descriptions $T$. The first type is single words - the 10k and 30k most common words in English. The second type is image descriptions - we prompt ChatGPT-3.5 to produce descriptions of images that include an object of a specific class. Repeating this process for all the ImageNet (IN) classes results in $\sim$28k unique image descriptions. We then evaluate the reconstruction of $r^{l}_{n}$ for different $m$’s and pools.
Effect of sparse set size $m$ and different pools. We experiment with $m\in\{4,8,16,32,64,128\}$ and the three text pools, and present the accuracy on 10% of ImageNet validation set in Figure 5. We approach the original classification accuracy with 128 text descriptions per neuron reconstruction $\hat{r}^{l}_{n}$. Using full descriptions outperforms using single words for the text pool, but the gap vanishes for larger $m$.
Qualitative results. We present the images with the largest second-order norms in Figure 6, and the corresponding top-4 text descriptions in Table 1. As shown, the found descriptions match the objects in the top 10 images. Moreover, some individual neurons correspond to multiple concepts (e.g. writing both toward “yacht” and a type of a car - “cabriolet”). This property is even more apparent if more nearest neighbors are presented (see Figure 14 for the top 50 nearest neighbors). This corroborates with previous literature on neurons’ polysemantic behavior (Elhage et al., 2022) - single neurons behave as a superposition of multiple interpretable features. This property will allow us to generate adversarial images in Section 5.1.
@@ 5.1 Automatic generation of adversarial examples
The sparse decomposition of $r_{n}^{l}$’s allows us to find overlapping concepts that neurons are writing to. We use these spurious cues to generate semantic adversarial images. Our pipeline, shown in Figure 1, mines for spurious words that correlate with the incorrect class in CLIP (e.g. “elephant”, that correlates with “dog”), combines them into image descriptions that include the correct class name (“cat”), and generates adversarial images by providing these descriptions to a text-to-image model. We explain the steps in the pipeline and provide quantitative and qualitative results.
@@
Experimental settings. We generate adversarial images for classifying between pairs of classes from CIFAR-10 (Krizhevsky, 2009). We use the 30k most common words as our pool $T$. We choose the top $100$ neurons from layers 8-10 for $\mathcal{N}$, and the top 25 words according to their contribution scores for prompting the LLM. We prompt LLaMA3 (Touvron et al., 2023) to generate 50 descriptions for each classification task (see prompt in Section A.7). We then filter out descriptions that include the class name and choose 10 random descriptions. We generate 10 images for each description with DeepFloyd IF text-to-image model (StabilityAI, 2023). This results in 100 images per experiment. We repeat the experiment 3 times and manually remove images that include $c_{2}$ objects or do not include $c_{1}$ objects.
We report three additional baselines. First, we repeat the same process with $100$ random neurons instead of the set $\mathcal{N}$. Second, we repeat the same generation process with sparse text decompositions computed from the first principle components of the indirect effects instead of the second-order effect. Third, we do not rely on the neuron decompositions, and instead prompt the language model with the words from $M$ for which their text representations are the most similar to $v$. Both for our pipeline and the baselines, we automatically filter out synonyms of $c_{2}$ from the phrases provided to the language model according to their sentence similarity to $c_{2}$ (Reimers & Gurevych, 2019).
Figure 7: Adversarial images generated by our method. For each binary classification task, we present the generated images, the input text to the text-to-image model (words from $W^{v}$ are bold), and an attribution map (Gandelsman et al., 2024) for the classification (areas that contribute to the incorrect class score are red). See additional results in Figure 16.
Quantitative results. The classification accuracy results for the adversarial images are presented in Table 2. The success rate of our adversarial images is significantly higher than the indirect effect baseline, the similar words baseline, and the random baseline, which succeeds only accidentally. For the task of generating “ship” images the will be missclassified as “truck”, no other baseline manged to generate any adversarial images, while ours generated 5.7 images on average.
Qualitative results. Figure 7 includes generated adversarial examples and the descriptions that were used in their generation. The presented attribution heatmaps (Gandelsman et al., 2024) show that the found spurious objects from $W_{v}$ contribute the most to the misclassification, while the object from the correct class (e.g. a horse in the left-most image) contributes the least. We provide more results for additional classification tasks (e.g. “stop-sign v.s. yield”) in Figure 16.
@@
horse $\rightarrow$ automobile 1.0 ($\pm$1.4) 2.8 ($\pm$3.7) 1.0 ($\pm$1.4) 5.3 ($\pm$1.9)
@@
ship $\rightarrow$ automobile 1.3 ($\pm$1.9) 0.0 ($\pm$0.0) 1.3 ($\pm$0.9) 7.0 ($\pm$4.5)
@@
Partial-LRP (Voita et al., 2019) 55.0 35.5 66.9
Rollout (Abnar & Zuidema, 2020) 61.8 42.6 74.0
LRP (Binder et al., 2016) 62.9 35.8 68.5
GradCAM (Selvaraju et al., 2017) 67.3 39.3 61.9
Chefer et al. (2021) 68.9 49.1 79.7
@@
Table 3: Segmentation performance on ImageNet-segmentation. Our zero-shot segmentation is more accurate than previous methods across all metrics.
Figure 8: Qualitative results on ImageNet-Segmentation (ViT-B-32). Our heatmaps capture more object parts than the first-order token decomposition of Gandelsman et al. (2024).
@@ 5.2 Zero-shot segmentation
Specifically, to generate a segmentation map for an image $I$, we find a set of neurons with the largest absolute value of the dot product with the encoded class name $c_{i}$ we aim to segment: $|\langle r_{n}^{l},M_{\text{text}}(c_{i})\rangle|$. We then average their spatial activation maps $p_{i}^{l,n}(I)$, standardize the average activations into $[0,1]$, and binarize the values into foreground/background segments by applying a threshold of 0.5.
Segmentation results. We provide results on ImageNet-Segmentation (Guillaumin et al., 2014), which includes foreground/background segmentation maps of ImageNet objects. We use activation maps from the top 200 neurons of layers 8-10. Table 3 presents a quantitative comparison to previous explainability methods. Our method outperforms other zero-shot segmentation methods across all standard evaluation metrics. We provide qualitative results before thresholding in Figure 8. While the first-order effects (“TextSpan”) highlight individual discriminative object parts, our heatmaps capture more parts of the full object.
@@ 6 Limitations and discussion
We analyzed the second-order effects of neurons on the CLIP representation and used our understanding to perform zero-shot segmentation and generate adversarial images. We present mechanisms that we did not analyze in our investigation and conclude with a discussion of broader impact and future directions.
@@
Future work and broader impact. The mass production of adversarial images can be harmful to systems that rely on neural networks (e.g., the adversarial attack that causes misclassification between “yield” and “stop sign” in Figure 16). Automatic extraction of such cases allows the defender to be prepared for them and, possibly, fine-tune the model on the generated images to avoid such attacks. We plan to investigate this approach to improve CLIP’s robustness in future work.
Currently, our attack pipeline relies on a few independent components, each of which has failure modes. For example, the language model can fail to generate a coherent sentence that includes many phrases from $W_{v}$, and can omit the class name $c_{2}$ or accidentally include the class name $c_{1}$. Additionally, the text-to-image model can fail to generate an image that follows the exact description and can drop crucial elements from the description. We believe that future improvements in the language and vision models will increase the success rate of our attack, and plan to continue to develop and improve it in the future.
@@ A.1 Second order ablations for ViT-L
We repeat the same experiments from Section 3.3 for ViT-L-14, trained on LAION dataset (Schuhmann et al., 2022). For this model, we only use 10% of ImageNet validation set. Here, the maximal drop in performance when ablating the second order is relatively smaller and is spread across more layers. Nevertheless, the same properties presented and discussed in Section 3.3 hold for this model.
@@ A.2 First order ablations
For the two models discussed above, ViT-B-32 and ViT-L-14, we provide the mean-ablation results for the first-order effects of MSA layers, as computed in Gandelsman et al. (2024). For each model, we present the performance before and after accumulative mean-ablation of all the first-order effects of MSA layers. As shown in Figure 13 and Figure 13, the neurons with the significant second-order effects appear right before the layers with the significant first-order effects.
@@ A.4 Additional sparse decomposition results
We provide additional examples of sparse decompositions of neurons in Table 4 and the images with the top norms for the second-order effects of the same neurons in Figure 15. As shown, the found descriptions match the objects in the top 10 images.
A.5 Concept discovery in images
We present an additional application - concept discovery in images. We aim to discover concepts in image $I$, by aggregating phrases that correspond to the neurons that are activated on $I$. Here, we start from the set of activated neurons $\mathcal{N}$ (for which $||\phi_{n}^{l}(I)||_{2}$ is above the 98th percentile of norms computed across ImageNet images). Similarly to the contribution score described in Section 5.1, we compute an image-contribution score $w_{j}^{I}$ for each phrase $j$ according to its combined weight in the decompositions of neurons in $\mathcal{N}$. Formally, $w_{j}^{I}$ is the overall sum of weights that each neuron in $\mathcal{N}$ assigns to $j$ in its decomposition, weighted by the neuron second-order norms: $w_{j}^{I}=\sum_{n\in{\mathcal{N}}}{\gamma_{j}^{l,n}||\phi_{n}^{l}(I)||_{2}}$. The phrases with the highest image-contribution score are picked to describe the image concepts.
Qualitative results. We present qualitative results for neurons and the top-10 discovered concepts from layer 9 of ViT-B-32 in Figure 9, when using the most common words as the pool. The number of neurons activated on these images, $|\mathcal{N}|$, is between 29 and 59, less than 2% of the neurons in the layer. Nevertheless, the top words extracted from these neurons relate semantically to the objects in the image and their locations. Surprisingly, the top word for each of the images appears only in one or two of the neuron sparse decompositions and is not spread across many activated neurons.
We acknowledge that while this application discovers meaningful concepts that correspond to the input images, there are other approaches for extract these concepts (e.g. sparsely decomposing the image representation, as shown in Bhalla et al. (2024)).
A.6 Derivations with Layer Normalization
@@
where $x\in\mathbb{R}^{d^{\prime}}$ is the input token, $\mu_{l},\sigma_{l}\in\mathbb{R}$ are the mean and standard deviation, and $\gamma,\beta\in\mathbb{R}^{d^{\prime}}$ are learned vectors. To include $A$ and $B$ in the second-order effect of a neuron flow, we replace the input-independent component in Equation 5, $PW_{VO}^{l^{\prime},h}w^{l^{\prime},n}$, with:
@@
where $LN^{l}$ is the normalization layer at layer $l$, that can be parameterized similarly to Equation 8 by $A^{l},B^{l}\in\mathbb{R}^{d^{\prime}}$. We modify the definition of the second-order effect accordingly:
@@
[…]
Figure 11: Mean-ablation of second order effects on ImageNet-R (ViT-B-32, layers 8-10). We repeat the evaluation in Figure 3 on ImageNet-R. The performance of different ablations follows the same trends as that of ImageNet.
@@
Figure 14: Images with largest second-order effect norm per neuron. We present the top images from 10% of ImageNet validation set, for the neurons in Table 1. Notice that additional concepts that are not captured by the top-4 descriptions in Table 1 are starting to appear.
Figure 15: Images with largest second-order effect norm per neuron. We present the top images from 10% of ImageNet validation set, for the neurons in Table 4.
@@
A.7 Prompts
We provide the prompt that was used for generating sentences given the set of words $W_{v}$, as presented in Section 5.1, in Table 5. This prompt is given to LLAMA3 model (Touvron et al., 2023).
Additionally, we provide the prompt that was used for generating the pool of ImageNet class descriptions, presented in Section 4. We prompt ChatGPT (GPT 3.5) with the prompt template provided in Table 6.
A.8 Compute